How much water must be evaporated from 240 gallons of a 3%salt solution to produce a 5% salt solution?
Answer by ptaylor(2198)

Let x=amount of water that needs to be evaporated
Now we know that the amount of pure salt in the 240 gal (0.03*240) has to equal the amount of pure salt in the solution after the evaporation takes place(0.05(240-x)). So our equation to solve is:
0.03*240=0.05*(240-x) get rid of parens and simplify
7.2=12-0.05x subtract 12 from each side
7.2-12=12-12-0.05x collect like terms
-4.8=-0.05x divide each side by 0.05
x=96 gal---------------------------------amount of water that needs to be evaporated
CK
0.03*240=0.05*144
7.2=7.2
Hope this helps---ptaylor